OpenStudy (anonymous):

If w = -5, what is the value of 3w2?

OpenStudy (anonymous):

3w^2?

hartnn (hartnn):

w^2 = (-5)^2= 25 so whats 3w^2 = 3*25 = ??

OpenStudy (anonymous):

75?

OpenStudy (anonymous):

yes

hartnn (hartnn):

75 it is :)